STUDY OF AIR-LIQUID INTERFACES USING SUM-FREQUENCY GENERATION SPECTROSCOPY: AP-PLICATIONS TO ATMOSPHERIC AEROSOLS

\underline{SANDHYA GOPALAKRISHNAN}, AND HEATHER C. ALLEN, Department of Chemistry, The Ohio State University, 100 W. 18th Ave., Columbus Ohio 43210.

The air-liquid interface of proxies for organic aerosols is investigated using broad bandwidth sum frequency generation (BBSFG) spectroscopy. The orientation of the fatty acid molecules was elucidated at the interface of a monolayer of oleic acid in water as well as in the pure liquid. This study is a prerequisite to the investigation of the heterogeneous reaction of ozone, a powerful atmospheric oxidant, with aerosols containing oleic acid. This reaction has not been directly investigated at the air-liquid interface and is of great interest in the context of the chemical transformation of aerosols containing organic molecules.